Solve :


x²y" -2xy' -4y=x²+2ln x

1
Expert's answer
2022-01-30T16:47:42-0500

Let us solve the differential equation "x^2y'' -2xy' -4y=x^2+2\ln x."

Let us use the transformation "x=e^t." Then "y'_x=y'_te^{-t},\ y''_{x^2}=(y''_{t^2}-y'_t)e^{-2t}."

We get the equation "e^{2t}(y''_{t^2}-y'_t)e^{-2t} -2e^ty'_te^{-t} -4y=e^{2t}+2t," which is equivalent to "y''_{t^2}-3y'_t -4y=e^{2t}+2t."

The characteristic equation "k^2-3k-4=0" of the differential equation "y''_{t^2}-3y'_t -4y=0" is equivalent to "(k+1)(k-4)=0," and hence has the roots "k_1=-1,\ k_2=4."

It follows that the general solution of the differential equation "y''_{t^2}-3y'_t -4y=e^{2t}+2t" is "y(t)=C_1e^{-t}+C_2e^{4t}+y_p," where "y_p=ae^{2t}+bt+d." Then "y_p'=2ae^{2t}+b,\ y_p''=4ae^{2t}."

It follows that

"4ae^{2t}-3(2ae^{2t}+b)-4(ae^{2t}+bt+d)=e^{2t}+2t," and hence

"-6ae^{2t}-4bt-3b-4d=e^{2t}+2t."

We conclude that "-6a=1,\ -4b=2,\ -3b-4d=0," and thus "a=-\frac{1}6,\ b=-\frac{1}2,\ d=-\frac{3}4b=\frac{3}8."

Therefore, the general solution of the differential equation "y''_{t^2}-3y'_t -4y=e^{2t}+2t" is "y(t)=C_1e^{-t}+C_2e^{4t}-\frac{1}6e^{2t}-\frac{1}2t+\frac{3}8."

We conclude that the general solution of the differential equation "x^2y'' -2xy' -4y=x^2+2\ln x" is "y(x)=C_1x^{-1}+C_2x^{4}-\frac{1}6x^{2}-\frac{1}2\ln x+\frac{3}8."